Evaluate the ∫1-siny dy/y+cosy


1
Expert's answer
2021-02-24T12:37:28-0500

Solution

"I=\int \frac{1-sin y}{y+cos y}dy"

Substitution x = y+cos y,  dx = 1-sin y:

"I=\int \frac{dx}{x}=ln|x|+C=ln|y+cos y|+C"

Answer

"\int \frac{1-sin y}{y+cos y}dy=ln|y+cos y|+C"